OpenStudy (anonymous):

csctheta=-5. what would sin, cos, tab, cot, and sec theta be?

sam (.sam.):

\[\csc \theta = \frac{1}{\sin \theta}\] ----------------------------------------- \[\frac{1}{\sin \theta}=-5\] \[\sin \theta=-\frac{1}{5}\] \[\theta = 191.53~,~348.47\] -------------------------------------------- For sin, cos, tab, cot, and sec, its a long step

OpenStudy (anonymous):

well, how would i solve for cos?

OpenStudy (anonymous):

sin^2(theta)+cos^2(theta)=1 1/25+cos^2(theta)=1 cos^2(theta)=24/25 cos theta=0.9797